OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y = 4 How do I solve this?

OpenStudy (haseeb96):

first convert the given question ( 2x+3x =4) into slope intercept form ( y=mx +b)

OpenStudy (haseeb96):

tell me what would be slope intercept form of this equation 2x + 3y = 4

OpenStudy (anonymous):

y=-2/3+4/3 Right?

OpenStudy (haseeb96):

@brimae97 tell me what is it slope intercept form

OpenStudy (cwrw238):

you omitted the x brimae

OpenStudy (haseeb96):

u forgot the x in your slope intercept form y = -2/3 x +4/3 this is your slope intercept form

OpenStudy (haseeb96):

now second step what is the slope in your slope intercept form for again review i write again y= mx +b here m is your slope so what would be the slope in this equation(y= -2/3 x + 4/3 )

OpenStudy (anonymous):

-2/3x is the slope

OpenStudy (haseeb96):

only -2/3 is your slope not -2/3x okay ?

OpenStudy (haseeb96):

so your slope is -2/3 okay

OpenStudy (haseeb96):

now if we again read the question statement ''The line through (1, -4) and parallel to 2x + 3y = 4'' it is saying that line is parallel to this equation so we know that when the lines are parallel then their slope will be equal m = m' here m = -2/3 so m'=-2/3 okay ? got it

OpenStudy (anonymous):

I got it, but how do you right the equation out in standard form? I don't understand that part.

OpenStudy (haseeb96):

i just take out it from question statement

OpenStudy (haseeb96):

now the point has given to u that is (1,-4) so (x',y')=(1,-4) okay and we have m' =-2/3 okay ?

OpenStudy (haseeb96):

now use the point slope form its formula is given as (y-y') = m' (x-x') since (x',y')=( 1,-4) m'=-2/3 just put these points in the above formula and solve it and show me what u get

OpenStudy (haseeb96):

y-(-4) = -2/3 (x-1) y+4=-2/3 (x-1) 3(y+4) =-2(x-1) 3y +12 =-2x +2 2x +3y +12 -2=0 2x +3y +10 =0 is your correct answer

OpenStudy (anonymous):

Thank you. I got half of it right, but I just didn't know how to put it in standard form, but now i do. Thank you so much!